OpenStudy (anonymous):

Which number can replace the variable to make these fractions equivalent? = A. 2 B. 3 C. 4 D. 5

OpenStudy (anonymous):

That would probably be A........?

OpenStudy (anonymous):

@leonardo0430

OpenStudy (leonardo0430):

i cant see the fraction >_> i have no account for k12

OpenStudy (jane11509):

I do

OpenStudy (leonardo0430):

wat are you doing here >_>

OpenStudy (anonymous):

@jane11509

OpenStudy (anonymous):

@jane11509

OpenStudy (jane11509):

I'm here o.o

OpenStudy (anonymous):

help me

OpenStudy (jane11509):

Ok

OpenStudy (jane11509):

Wat is the variable I cannot see it o.o

OpenStudy (anonymous):

n/4= 15/20

OpenStudy (anonymous):

@jane11509

OpenStudy (jane11509):

Okay let me see

OpenStudy (anonymous):

so

OpenStudy (leonardo0430):

hold on let me check

OpenStudy (leonardo0430):

can you type in the fraction

OpenStudy (anonymous):

n/4 = 15/20

OpenStudy (jane11509):

It's B 3/4 = 15/20

OpenStudy (leonardo0430):

its 2 so it has to be A

OpenStudy (anonymous):

a

OpenStudy (leonardo0430):

:p what did i say

OpenStudy (anonymous):

just making sure